# Online Combinatorial Optimization with Graphical Dependencies

### v1

@@ Abstract
Most existing work in online stochastic combinatorial optimization assumes that inputs are drawn from independent distributions—a strong assumption that often fails in practice. At the other extreme, arbitrary correlations are equivalent to worst-case inputs via Yao’s minimax principle, making good algorithms often impossible. This motivates the study of intermediate models that capture mild correlations while still permitting non-trivial algorithms.
@@ 1 Introduction
To overcome worst-case inapproximability, much of the recent work in online combinatorial optimization has focused on stochastic input models. A central example is the prophet model, where each request $i\in[n]$ is independently drawn from a known distribution $\mathcal{D}_{i}$, and the algorithm aims to compete with the expected offline optimum over the product distribution $\mathcal{D}=\mathcal{D}_{1}\times\dots\times\mathcal{D}_{n}$. This model has led to $O(1)$-competitive algorithms for a wide range of problems, including online matchings CMS (10); GW (19); EFGT (20), combinatorial auctions FGL (15); DFKL (20); CC (23), Facility Location and Steiner Tree GGLS (08); AFGS (22).
However, these guarantees critically rely on the independence assumption. In the presence of arbitrary correlations between inputs, such guarantees collapse. This failure is formalized by Yao’s minimax principle Yao (77), which shows that online algorithms for known correlated distributions are no easier than those for worst-case adversarial inputs. Since real-world data often exhibits non adversarial dependencies—and arbitrary correlations lead back to worst-case hardness—this motivates the study of online stochastic optimization under mild, structured correlations.
@@ 1.1 Our Results and High-Level Techniques
In the $p$-sample model, a set of requests is chosen adversarially but then a random $p$-fraction is independently revealed in advance to the algorithm as sample, while the remaining $(1-p)$ fraction appear one by one in an adversarial order (Definition 3.4). Most algorithms in this model are also monotone—their cost only decreases if the sample set increases. We show that any monotone algorithm in the $p$-sample model implies an algorithm in the MRF model, which enables us to leverage existing $p$-sample algorithms.
@@
As a concrete application, we obtain $O(\Delta)$-competitive algorithms for Facility Location and Steiner Tree, using prior $p$-sample algorithms or their small modifications AFGS (22). Notably, our result provides the first nontrivial approximation guarantees for these problems in the presence of structured correlations.
@@
For online Facility Location and online Steiner Tree in the MRF arrival model with maximum weighted degree $\Delta$, there exist $O(\Delta)$-competitive algorithms.
@@ 2 Preliminaries on Markov Random Fields
For our applications, we will be interested in distributions over non-negative reals, i.e., $\Omega=\mathbb{R}_{\geq 0}^{n}$. To build some intuition on MRFs, notice that in the absence of edges, they are equivalent to product distributions. The general MRF model captures several important graphical models as special cases:
  • Ising model: This corresponds to having a simple graph (i.e., no edge with $3$ or more vertices) with $\Omega=\{-1,1\}^{n}$ and the function $\psi_{e}$ taking value either $+J_{e}$ or $-J_{e}$ for some $J_{e}\in\mathbb{R}$, depending on whether the two end points agree or disagree with each other.
  • (Anti-)Ferromagnetic Ising model: Special case of Ising model with identical edge weights, i.e., $J_{e}=J$. It’s called ferromagnetic or anti-ferromagnetic depending on the sign of $J$ and its magnitude is called the inverse temperature.
  • Sherrington–Kirkpatrick model: Special case of Ising model where the underlying graph is complete and the edge weights $\psi_{e}$ are chosen from Gaussian distribution.
  • Potts model: When we have a simple graph but general $\Omega$.
@@
For any MRF $\mathcal{M}=(\Omega,E,\{\psi_{i}\}_{i\in[n]},\{\psi_{e}\}_{e\in E})$, the weighted maximium degree
@@ 3.1 Model: Subadditive Coverage Problems
Similarly to GGLS (08), let $V$ denote a set of demand points and $E$ denote the ground set of elements. For any set of demand points $D\subseteq V$, there is a collection of feasible solutions denoted by $\mathcal{F}_{D}\subseteq 2^{E}$. First, we formally define a general coverage problem, which roughly requires two properties: (1) we must be able to extend any solution for a set of demands to a solution for a superset of the demands, and (2) the combination of two feasible solutions for two different demand sets should be feasible for their union.
Definition 3.1 (Coverage constraints AFGS (22)).
A collection of feasible solutions $\{\mathcal{F}_{D}\}_{D\subseteq V}$ such that
@@
Now we can define a subadditive coverage problem.
Definition 3.2 ((Online) Subadditive Coverage problem).
Let $V$ denote a set of demand points and $E$ denote the ground set of elements. We are also given a collection of generalized coverage constraints $\{\mathcal{F}_{D}\}_{D\subseteq V}$. Any subset of elements $S\subseteq E$ can be selected (bought) by paying a monotone subadditive [^1] For all $S_{1},S_{2}\subseteq E$, we have $c(S_{1})+c(S_{2})\geq c(S_{1}\cup S_{2})$. cost $c(S)\in\mathbb{R}_{\geq 0}$. For a set of demands $D\subseteq V$, the goal is to select a subset of elements $S\in\mathcal{F}_{D}$ while minimizing cost $c(S)$.
In the online setting, a sequence of demands $D=\{v_{1},\dots,v_{n}\}$ arrive one by one (possibly from some known correlated probability distribution). Starting with $F_{0}=\varnothing$, after each arrival $v_{i}$, we must pick a subset of elements $S_{i}\subseteq E$ to select such that $F_{i}=F_{i-1}\cup S_{i}$ satisfies $F_{i}\in\mathcal{F}_{D_{\leq i}}$, where $D_{\leq i}:=\{v_{1},\dots,v_{i}\}$ denotes the set of all current demand points. Our objective is to choose a feasible solution to minimize $c(F_{n})$.
@@
  • Steiner Tree: We have a graph $G=(V,E)$ with a fixed root vertex $r\in V$. Demands are the vertices of the graph and elements are the edges. For a set of demands $D\subseteq V$, a feasible solution $F\in\mathcal{F}_{D}$ is a set of edges such that all demands are connected to the root in the graph $(V,F)$. Each edge has an associated cost $c(e)$, and costs are additive. The classic result of IW (91) gives a tight $\Theta(\log n)$-competitive algorithm for online adversarial arrivals.
  • Facility Location: The demand points $v$ (clients) are points in a metric space $M$. Elements include both points $f\in M$ (facilities) and connections $(f,v)\in M\times M$ between a facility and a client. For a set of clients $D\subseteq M$, a feasible solution is a set of open facilities and connections such that each client is connected to an open facility. Costs are additive: each facility $f\in M$ has an opening cost $c(f)$ and each connection $(f,v)$ has an assignment cost $c((f,v))=d_{M}(f,v)$ given by the metric distance. We know tight $\Theta(\log n/\log\!\log n)$-competitive algorithm for online adversarial arrivals Fot (08); Mey (01).
[…]
@@ 3.2 Main Result and Proof Overview
This section presents our reduction for online Subadditive Coverage from the MRF arrival model to a model without correlations where some of the input demands are given upfront as samples. We call this latter setting the $p$-sample model. Using known results in this framework, our reduction would imply approximation guarantees in the MRF setting. Before presenting our main theorem, we formally define these two models.
@@
In the single-sample $\Delta$-MRF prophet model, the setup is the same as $\Delta$-MRF prophet model, except that the underlying MRF distribution is unknown to the algorithm, and instead it is provided a single $n$-dimensional sample from this MRF distribution.
@@
There is an adversarially chosen unknown value set $V=\{v_{1},v_{2},\dots,v_{n}\}$. The algorithm is given a random sample $S\subseteq V$ before the start such that each $v_{i}$ belongs to $S$ independently with a known probability $p\in[0,1]$. The remaining values $R=V\setminus S$ are the real values, forming the demand set $D=R$ and arriving in an adversarial order.
@@
For any $p\in[0,1]$, if there exists an $\alpha(p)$-competitive monotone algorithm for a subadditive coverage problem in the $p$-sample independent model, then there exists an $\alpha\left(\frac{1}{2}e^{-8\Delta}\right)$-competitive algorithm for the same problem in the $\Delta$-MRF model. Specifically, an $O\left(\log({1}/{p})\right)$-approximation in the $p$-sample model yields an $O(\Delta)$-approximation in the MRF model. Moreover, this reduction holds even from the single-sample $\Delta$-MRF prophet model to the $p$-sample independent model.
@@
An adversary chooses arbitrary values $v^{1}_{1},\dots,v^{1}_{n},v^{-1}_{1},\dots,v^{-1}_{n}$ unknown to the algorithm. We then sample $\bm{\sigma}\in\{-1,1\}^{n}$ from an unknown symmetric [^2] Symmetric means that for any $\bm{\sigma}\in\{-1,1\}^{n}$, both $\bm{\sigma}$ and $-\bm{\sigma}$ have the same probability. $\Delta$-MRF distribution $\mathcal{D}_{\sigma}$ over $\{-1,1\}^{n}$. When a sample $\bm{\sigma}\sim\mathcal{D}_{\sigma}$ is drawn, the values $\bm{v}^{\bm{\sigma}}=(v^{\sigma_{1}}_{1},\dots,v^{\sigma_{n}}_{n})$ become the sample set $S=\bm{v}^{\bm{\sigma}}$ and are revealed to the algorithm offline, while the values $\bm{v}^{-\bm{\sigma}}=(v^{-\sigma_{1}}_{1},\dots,v^{-\sigma_{n}}_{n})$ are the real values $R=\bm{v}^{-\bm{\sigma}}$ and arrive sequentially as demands $D=R$. (Note that symmetry implies $S$ and $R$ are identically distributed.)
@@
We prove Theorem 3.5 by sequentially reducing from the initial single-sample MRF prophet problem to the Game of Googol MRF, to the $(\frac{1}{2},p)$-sample, and finally to the $p$-sample independent model, and showing how we can preserve the approximation ratios. The sequence of reductions is shown in Figure 1, and the main lemmas are shown in Sections 3.3.1, 3.3.2, and 3.3.3.
@@ 3.3.1 Reducing Single-Sample $\Delta$-MRF Prophet to Game of Googol MRF
We show that for any realization of $V$, the distribution $\mathcal{D}^{V}_{\sigma}$ of $\sigma$ conditional on $V$ can be represented as a $2\Delta$-MRF. (Note that this distribution depends on $V$, so is unknown to the algorithm, but we only need its existence.) Let $\psi_{i}$ for $i\in[n]$ and $\psi_{e}$ for $e\in E$ be the vertex and edge potentials in the MRF $\mathcal{D}$ respectively. From Definition 2.1, we know that for any fixed $v_{1},\dots,v_{n}$,
@@
Using this, we compute the probability mass function of $\mathcal{D}^{V}_{\sigma}$. We let $\bm{v}^{\bm{\sigma}}:=(v_{1}^{\sigma_{1}},\dots,v_{n}^{\sigma_{n}})$, and likewise define $\bm{v}^{-\bm{\sigma}}$. Then,
$\Pr_{\sigma^{\prime}\sim\mathcal{D}^{V}_{\sigma}}[\sigma^{\prime}=\sigma]$ $\propto\Pr_{S\sim\mathcal{D}}[S=\bm{v}^{\bm{\sigma}}]\cdot\Pr_{R\sim\mathcal{D}}[R=\bm{v}^{-\bm{\sigma}}]$
@@
Therefore, $\mathcal{D}^{V}_{\sigma}$ can be represented as an MRF with potential functions $\hat{\psi}_{i}(\sigma_{i})=0$ and $\hat{\psi}_{e}(\sigma_{e})=\psi_{e}(v^{\sigma}_{e})+\psi_{e}(v^{-\sigma}_{e})$, and observe that by our construction the maximum degree of this MRF is at most $2\Delta$. Additionally, it is not hard to see that $\mathcal{D}_{\sigma}^{V}$ is symmetric, has uniform marginals, and is known to the algorithm as it is determined by $V$ and $\mathcal{D}$. Hence, we this gives a valid game of googol instance.
Completing the proof, the original single-sample prophet instance can be generated by first sampling $V$ and $\sigma\sim\mathcal{D}^{V}_{\sigma}$, and then setting $S=\bm{v}^{\bm{\sigma}}$ and $R=\bm{v}^{-\bm{\sigma}}$. With these sets, we can run $\alpha$-competitive algorithm for the $(2\Delta)$-MRF Game of Google model to get an $\alpha$-competitive ratio in the $\Delta$-MRF setting. ∎

### v2

@@ Abstract
Most existing work in online stochastic combinatorial optimization assumes that inputs are drawn from independent distributions—a strong assumption that often fails in practice. At the other extreme, arbitrary correlations are equivalent to worst-case inputs via Yao’s minimax principle, making good algorithms often impossible. This motivates the study of intermediate models that capture mild correlations while still permitting nontrivial algorithms.
@@ 1 Introduction
To overcome worst-case inapproximability, much recent work in online combinatorial optimization has focused on stochastic input models. A central example is the prophet model, where each request $i\in[n]$ is independently drawn from a known distribution $\mathcal{D}_{i}$, and the algorithm aims to compete with the expected offline optimum over the product distribution $\mathcal{D}=\mathcal{D}_{1}\times\dots\times\mathcal{D}_{n}$. This model has led to $O(1)$-competitive algorithms for a wide range of problems, including online matchings CMS (10); GW (19); EFGT (20), combinatorial auctions FGL (15); DFKL (20); CC (23), Facility Location and Steiner Tree GGLS (08); AFGS (22).
However, these guarantees critically rely on the independence assumption. In the presence of arbitrary correlations between inputs, such guarantees collapse. This failure is formalized by Yao’s minimax principle Yao (77), which shows that online algorithms for known correlated distributions are no easier than those for worst-case adversarial inputs. Since real-world data often exhibits non-adversarial dependencies—and arbitrary correlations lead back to worst-case hardness—this motivates the study of online stochastic optimization under mild, structured correlations.
@@ 1.1 Our Results and High-Level Techniques
In the $p$-sample model, a set of requests is chosen adversarially but then a random $p$-fraction is independently revealed in advance to the algorithm as sample, while the remaining (1p)(1-p) fraction appear one by one in an adversarial order (Definition 3.4). Most algorithms in this model are also monotone—their cost only decreases if the sample set increases. We show that any monotone algorithm in the $p$-sample model implies an algorithm in the MRF model, which enables us to leverage existing $p$-sample algorithms.
@@
As a concrete application, we obtain $O(\Delta)$-competitive algorithms for Facility Location and Steiner Tree, using prior $p$-sample algorithms or their small modifications AFGS (22). Notably, our result provides the first nontrivial approximation guarantees for these problems in the presence of structured correlations. Moreover, it is particularly interesting that we only require a single-sample from the underlying distribution, since learning MRF distributions is a challenging problem and an active area of research. Our theorem permits one to bypass learning the correlated MRF distribution to design an algorithm directly.
@@
For online Facility Location and online Steiner Tree in the MRF arrival model with maximum weighted degree $\Delta$, there exist $O(\Delta)$-competitive algorithms given only a single $n$ dimensional sample from the underlying MRF.
@@ 2 Preliminaries on Markov Random Fields
To build some intuition on MRFs, notice that in the absence of edges, they are equivalent to product distributions. The general MRF model captures several important graphical models as special cases:
  • Ising model: This corresponds to having a simple graph (i.e., where no edge has $3$ or more vertices) with $\Omega=\{-1,1\}^{n}$ and the function $\psi_{e}$ taking value either $J_{e}$ or $J_{e}$ for some $J_{e}\in\mathbb{R}$, depending on whether the two end points agree or disagree with each other.
  • (Anti-)Ferromagnetic Ising model: This is a special case of Ising model with identical edge weights, i.e., $J_{e}=J$. It’s called ferromagnetic or anti-ferromagnetic depending on the sign of $J$, and its magnitude is called the inverse temperature.
  • Sherrington–Kirkpatrick model: This is another special case of Ising model where the underlying graph is complete and the edge weights $\psi_{e}$ are chosen from Gaussian distribution.
  • Potts model: This is where we have a simple graph, and each vertex has $q$ states $\Omega_{i}=\{1,\dots,q\}$.
@@
For any MRF $\mathcal{M}=(\Omega,E,\{\psi_{i}\}_{i\in[n]},\{\psi_{e}\}_{e\in E})$, the weighted maximum degree
@@ 3.1 Model: Subadditive Coverage Problems
Similarly to GGLS (08), let $V$ denote a ground set of possible demand points and $E$ denote the ground set of elements used to construct a solution. For any set of demand points $D\subseteq V$, we have a collection of feasible solutions denoted by $\mathcal{F}_{D}\subseteq 2^{E}$. To capture general coverage constraints, these solutions sets should have two key properties: (1) we must be able to extend any solution for a set of demands to a solution for a superset of the demands, and (2) the union of two feasible solutions for two different demand sets should be a feasible solution for the union of demand sets. We formalize this notion of coverage constraints bellow.
Definition 3.1 (Generalized coverage constraints AFGS (22)).
Given a universe of demands $V$ and a universe of solution elements $E$, a family of generalized coverage constraints is a collection of feasible solutions $\mathcal{F}_{D}\subseteq 2^{E}$ for each $D\subseteq 2^{V}$ such that
@@
Using this notion of general coverage constraints, we can define a subadditive coverage problem as follows.
Definition 3.2 ((Online) Subadditive Coverage).
In a subadditive coverage problem, we have a universe of demand points $V$, a universe of solution elements $E$, a collection of generalized coverage constraints $\{\mathcal{F}_{D}\}_{D\subseteq V}$, and a monotone subadditive [^1] A function $c$ is subadditive if, for all $S_{1},S_{2}\subseteq E$, we have $c(S_{1})+c(S_{2})\geq c(S_{1}\cup S_{2})$. cost function $c:2^{E}\to\mathbb{R}_{\geq 0}$. Any subset of elements $S\subseteq E$ can be selected (bought) by paying cost $c(S)\in\mathbb{R}_{\geq 0}$. Given a set of demands $D\subseteq V$, the goal is to select a subset of elements $F\in\mathcal{F}_{D}$ while minimizing cost $c(F)$.
In the online setting, a sequence of demands $D=\{v_{1},\dots,v_{n}\}$ arrive one by one (possibly from some known arrival model). Starting with $F_{0}=\varnothing$, after each arrival $v_{i}$, we must select a subset of new elements $S_{i}\subseteq E$ such that the total set of elements selected $F_{i}=F_{i-1}\cup S_{i}$ satisfies $F_{i}\in\mathcal{F}_{D_{\leq i}}$, where $D_{\leq i}:=\{v_{1},\dots,v_{i}\}$ denotes the set of all current demand points. Our objective is to choose a feasible solution to minimize $c(F_{n})$.
@@
  • Steiner Tree: We have a graph $G=(V,E)$ with a fixed root vertex $r\in V$. Demands are the vertices of the graph and solution elements are the edges. For a set of demands $D\subseteq V$, a feasible solution $F\in\mathcal{F}_{D}$ is a set of edges such that all demands are connected to the root in the graph $(V,F)$. Each edge has an associated cost $c(e)$, and costs are additive. The classic result of IW (91) gives a tight $\Theta(\log n)$-competitive algorithm for online adversarial arrivals.
  • Facility Location: The demand points $v$ (clients) are points in a metric space $M$. Solution elements include both points $f\in M$ (facilities) and connections $(f,v)\in M\times M$ between a facility and a client. For a set of clients $D\subseteq M$, a feasible solution is a set of open facilities and connections such that each client is connected to an open facility. Costs are additive: each facility $f\in M$ has an opening cost $c(f)$ and each connection $(f,v)$ has an assignment cost $c((f,v))=d_{M}(f,v)$ given by the metric distance. We know tight $\Theta(\log n/\log\!\log n)$-competitive algorithm for online adversarial arrivals Fot (08); Mey (01).
  • Set Cover: Elements $E=[m]$ consist of indices of sets $S_{1},\dots,S_{m}$, and demands consist of points $j\in\bigcup_{i}S_{i}$. For a set of demands $D$, a feasible solution $F\in\mathcal{F}_{D}$ is a set on indices such that $D\subseteq\bigcup_{i\in F}S_{i}$. Each set $S_{i}$ has an associated cost $c(i)$, and costs are additive. Unless $\text{NP}\subseteq\text{BPP}$, we know $\Theta(\log m\cdot\log n)$ is the tight competitive ratio for online adversarial arrivals AAA (03); Kor (04).
@@ 3.2 Main Result and Proof Overview
This section presents our reduction for online subadditive coverage from the MRF arrival model (i.e. where $D$ is sampled from a known MRF distribution) to a model without correlations where some of the input demands are given upfront as samples. We call this latter setting the $p$-sample model. Using known results in this framework, our reduction would imply approximation guarantees in the MRF setting. Before presenting our main theorem, we formally define these two models.
@@
In the single-sample $\Delta$-MRF prophet model, the setup is the same as $\Delta$-MRF prophet model, except that the underlying MRF distribution is unknown, and instead the algorithm is provided a single $n$-dimensional sample from this MRF distribution.
@@
There is an unknown set of possible demands $V=\{v_{1},v_{2},\dots,v_{n}\}$. The algorithm is given a random sample $S\subseteq V$ before the start such that each $v_{i}$ belongs to $S$ independently with a known probability $p\in[0,1]$. The remaining values $R=V\setminus S$ are the real values, forming the demand set $D=R$ and arriving in an adversarial order.
@@
For any $p\in[0,1]$, if there exists an $\alpha(p)$-competitive monotone algorithm for a subadditive coverage problem in the $p$-sample independent model, then there exists a $2\alpha\left(\frac{1}{2}e^{-8\Delta}\right)$-competitive algorithm for the same problem in the $\Delta$-MRF model. Specifically, an $O\left(\log({1}/{p})\right)$-approximation in the $p$-sample model yields an $O(\Delta)$-approximation in the MRF model. Moreover, this reduction holds even from the single-sample $\Delta$-MRF prophet model to the $p$-sample independent model.
@@
An adversary chooses arbitrary values $v^{1}_{1},\dots,v^{1}_{n},v^{-1}_{1},\dots,v^{-1}_{n}$ unknown to the algorithm. We then sample $\bm{\sigma}\in\{-1,1\}^{n}$ from an unknown symmetric [^2] Symmetric means that for any $\bm{\sigma}\in\{-1,1\}^{n}$, both $\bm{\sigma}$ and $\bm{\sigma}$ have the same probability. $\Delta$-MRF distribution $\mathcal{D}_{\sigma}$ over $\{-1,1\}^{n}$. When a sample $\bm{\sigma}\sim\mathcal{D}_{\sigma}$ is drawn, the values $\bm{v}^{\bm{\sigma}}=(v^{\sigma_{1}}_{1},\dots,v^{\sigma_{n}}_{n})$ become the sample set $S=\bm{v}^{\bm{\sigma}}$ and are revealed to the algorithm offline, while the values $\bm{v}^{-\bm{\sigma}}=(v^{-\sigma_{1}}_{1},\dots,v^{-\sigma_{n}}_{n})$ are the real values $R=\bm{v}^{-\bm{\sigma}}$ and arrive sequentially as demands $D=R$. (Note that symmetry implies $S$ and $R$ are identically distributed.)
@@
We prove Theorem 3.5 by sequentially reducing from the initial single-sample MRF prophet problem to the Game of Googol MRF, to the $(\frac{1}{2},p)$-sample, and finally to the $p$-sample independent model. Figure 1 shows this reduction, and the main lemmas are given in Sections 3.3.1, 3.3.2, and 3.3.3.
@@ 3.3.1 Reducing Single-Sample $\Delta$-MRF Prophet to Game of Googol MRF
We show that for any realization of $V$, the distribution $\mathcal{D}^{V}_{\bm{\sigma}}$ of $\bm{\sigma}$ conditional on $V$ can be represented as a $2\Delta$-MRF. (Note that this distribution depends on $V$, so is unknown to the algorithm, but we only need its existence.) Let $\psi_{i}$ for $i\in[n]$ and $\psi_{e}$ for $e\in E$ be the vertex and edge potentials in the MRF $\mathcal{D}$ respectively. From Definition 2.1, we know that for any fixed $v_{1},\dots,v_{n}$,
@@
Using this, we compute the probability mass function of $\mathcal{D}^{V}_{\bm{\sigma}}$. The idea is that $V=V(S,R,\bm{\sigma}^{\prime})$ depends on three independent sources of randomness: the MRF samples $S\sim\mathcal{D}$ and $R\sim\mathcal{D}$, as well as the uniform $\bm{\sigma}^{\prime}\sim\mathrm{Unif}\{-1,1\}^{n}$. We show that, if given only the result $V$, then the conditional distribution on $\bm{\sigma}^{\prime}$ can be expressed in terms of the distributions of $S$ and $R$ using Bayes’ theorem:
$\Pr_{\bm{\sigma}^{\prime},S,R}[\bm{\sigma}^{\prime}=\bm{\sigma}\mid V(S,R,\bm{\sigma}^{\prime})=V]$ $=\frac{\Pr_{\bm{\sigma}^{\prime},S,R}[V(S,R,\bm{\sigma}^{\prime})=V\mid\bm{\sigma}^{\prime}=\bm{\sigma}]\cdot\Pr_{\bm{\sigma}^{\prime}}[\bm{\sigma}^{\prime}=\bm{\sigma}]}{\Pr_{\bm{\sigma}^{\prime},S,R}[V(S,R,\bm{\sigma}^{\prime})=V]}$
$=\frac{\Pr_{\bm{\sigma}^{\prime},S,R}[V(S,R,\bm{\sigma}^{\prime})=V\mid\bm{\sigma}^{\prime}=\bm{\sigma}]\cdot 2^{-n}}{\Pr_{\bm{\sigma}^{\prime},S,R}[V(S,R,\bm{\sigma}^{\prime})=V]}$
$\propto\Pr_{\bm{\sigma}^{\prime},S,R}[V(S,R,\bm{\sigma}^{\prime})=V\mid\bm{\sigma}^{\prime}=\bm{\sigma}]$
$=\Pr_{S,R}[V(S,R,\bm{\sigma})=V]$
$=\Pr_{S}[S=v^{\bm{\sigma}}]\cdot\Pr_{R}[R=v^{-\bm{\sigma}}]$
Next, let $\bm{v}^{\bm{\sigma}}:=(v_{1}^{\sigma_{1}},\dots,v_{n}^{\sigma_{n}})$, and likewise define $\bm{v}^{-\bm{\sigma}}$. We have
$\Pr_{\bm{\sigma^{\prime}}\sim\mathcal{D}^{V}_{\bm{\sigma}}}[\bm{\sigma^{\prime}}=\bm{\sigma}]$ $\propto\Pr_{S\sim\mathcal{D}}[S=\bm{v}^{\bm{\sigma}}]\cdot\Pr_{R\sim\mathcal{D}}[R=\bm{v}^{-\bm{\sigma}}]$
@@
Therefore, $\mathcal{D}^{V}_{\bm{\sigma}}$ can be represented as an MRF with potential functions $\hat{\psi}_{i}(\sigma_{i})=0$ and $\hat{\psi}_{e}(\bm{\sigma}_{e})=\psi_{e}(v^{\sigma}_{e})+\psi_{e}(v^{-\sigma}_{e})$, and observe that by our construction the maximum degree of this MRF is at most $2\Delta$. Additionally, it is not hard to see that $\mathcal{D}_{\bm{\sigma}}^{V}$ is symmetric, has uniform marginals, and is known to the algorithm as it is determined by $V$ and $\mathcal{D}$. Hence, we this gives a valid game of googol instance.
Completing the proof, the original single-sample prophet instance can be generated by first sampling $V$ and $\bm{\sigma}\sim\mathcal{D}^{V}_{\bm{\sigma}}$, and then setting $S=\bm{v}^{\bm{\sigma}}$ and $R=\bm{v}^{-\bm{\sigma}}$. With these sets, we can run $\alpha$-competitive algorithm for the $(2\Delta)$-MRF Game of Google model to get an $\alpha$-competitive ratio in the $\Delta$-MRF setting. ∎